OpenStudy (anonymous):

find the least common multiple of 8v^4u^8w^5 and 6u^6w^3 The ^means to the power of the next number PLEASE HELP!!!

OpenStudy (zzr0ck3r):

2^3*v^4*u^8*w^5 3*2*u^6*w^3 we take the highest power of each base 2^3*3*u^8*v^4*w^5 24u^8v^4w^5

OpenStudy (anonymous):

im confused where the 24 came from

OpenStudy (zzr0ck3r):

2^3*3 = 8*3 = 24

OpenStudy (zzr0ck3r):

the highest power of 2 is 3 the highest power of 3 is 1 2^3*3^1 = 24

OpenStudy (anonymous):

ok thank you

OpenStudy (zzr0ck3r):

np

OpenStudy (zzr0ck3r):

what is the least common multiply of 120 and 90 same thing 120 = 3*5*2*2*2 90 = 3*^2*5*2 highest power of 2 is 2^3 highest power of 3 is 3^2 highest power of 5 is 5^2 so lcm = 2^3*3^2*5^2 = 8*9*25 = 1800

OpenStudy (anonymous):

are you good with functions? there is a question i have been stuck on and you seem like you know quite a bit.

OpenStudy (zzr0ck3r):

sure